# Synthesizing Programmatic Reinforcement Learning Policies with Large Language Model Guided Search

### v1

@@ 4.3 Optimizing the program policy with Scheduled Hill Climbing
Although the aforementioned techniques can already instruct the LLM to produce DSL programs that solve the target task to a certain degree, the episodic return can still be far from optimal. Since best-in-class LLMs are all proprietary APIs with no fine-tuning service; therefore, conventional gradient-based policy optimization is out of the question. To further optimize the program generated by the LLM, we explore initializing search populations of search algorithms using LLM-generated programs. Via extensive experiments with the HC, CEM, and CEBS search algorithms on both program space and latent space whenever applicable, we discover a key improvement that is crucial when initializing search from LLM-generated programs – program-environment interaction scheduling for a more efficient allocation of the interaction budget. The intuition is that LLMs can often provide good initialization and therefore it makes sense to keep a small search budget, i.e. with a smaller population. After a while, if an optimal program is not yet found, we can gradually increase the budget and facilitate broader exploration. We design a scheduler based on this intuition:
@@
where $n$ represents the number of evaluated programs, $k(n)$ denotes a function indicating the current number of neighborhood programs based on the number of evaluated programs $n$. This function’s logarithm is a linear combination of the logarithms of two hyperparameters, $K_{start}$ and $K_{end}$, which signify the initial and terminal numbers of neighbors to search, respectively. The variable $r(n)$ governs the linear ratio, which itself is a sinusoidal function that gradually increases from 00 to $1$ throughout the evaluation of a total of $N$ programs (see Appendix A). With this scheduler, the number of neighbors $k$ is a function of the execution budget used, growing from $K_{start}$ to $K_{end}$. We apply this scheduler to Hill Climbing and call this method Scheduled Hill Climbing (Scheduled HC).
Overall, our best recipe consists of the following steps. First, we sample DSL programs for the LLM using the aforementioned techniques. Next, these programs are evaluated in the environment for the episodic return. In some easy tasks, some generated programs might have achieved perfect returns and thus are optimal programs. If no optimal program is found, we sort the programs based on the total reward decreasingly. Next, programs are selected from the sorted list, each serves as an initial program for the HC search, which is the best-performing search algorithm from our extensive experiments. HC searches $k$ neighbors at each step following the scheduler in Eq. 1. This process is repeated until either the optimal program is found or we meet the maximum program-environment interaction budget allowed ($N$). Figure 3 presents an overview of our framework.
@@
Figure 4: Efficiency in the Karel domain. We compare our proposed LLM-guided search (LLM-GS) framework against existing methods, LEAPS, HPRL, CEBS, and HC in the Karel and Karel-Hard problem sets. The results show that our LLM-GS is significantly more efficient than these methods.
@@ 5.1 Evaluation setup
PRL tasks. We evaluate our proposed framework LLM-guided search (LLM-GS) using the Karel tasks from the two problem sets: Karel [57] (StairClimber, Maze, FourCorners, TopOff, Harvester, and CleanHouse) and Karel-Hard [37] (DoorKey, OneStroke, Seeder, and Snake). More details of the task can be found in Appendix C.
Baselines. We compare our proposed framework with the following methods: LEAPS [57], HPRL [37], CEBS [10], and Hill Climbing [10] (the current state-of-the-art). More details of these methods are presented in Appendix D.
Metrics. For a fair comparison of the sample efficiency, we use the number of programs evaluated to represent the program-environment interaction budget. A method is a more sample-efficient PRL algorithm if it achieves a higher average episodic return on a fixed budget. Specifically, for each task, the number of task variances is $C$. One program evaluated means a program is executed on all $C$ task variances to obtain the average return. We set the number of task variances $C=32$, the maximum number of program evaluation $N=10^{6}$, i.e., the interaction budget.
Setup. We use gpt-4-turbo [1] (temperature=1.0, top_p=1.0) as our LLM module to generate the initial search population with 32 programs. The scheduler of our proposed Scheduled HC starts from $K_{start}=32$ to $K_{end}=2048$. We evaluate our LLM-GS and HC with $32$ random seeds and $5$ seeds for LEAPS, HPRL, and CEBS.
@@ 5.2 LLM-guided search significantly improves the sample efficiency
Figure 4 presents our full experimental result, indicating that our proposed LLM-GS surpasses all the existing methods by a large margin on almost all the tasks. On the Karel set, our framework completely solves these tasks and exhibits the best sample efficiency. Also, LLM-GS performs extraordinarily well on Seeder, OneStroke, and DoorKey, which are among the hardest of the Karel-Hard set. We highlight our DoorKey result, with an efficiency improvement from not even converging at 1M to converge within around 50K. Existing PRL algorithms struggle at DoorKey due to the fact that it is a two-stage task with a sparse reward function. Therefore, most of the baselines cannot escape the local maxima and converge around $0.5$. Curiously, our method starts at an episodic return of $\approx 0.5$ from the beginning and converges quickly without being trapped by the local maxima. To analyze, we examine one output program from LLM shown in Figure 5. The LLM clearly understands that DoorKey is a two-stage task and thus generates a corresponding two-stage structured program, which is much easier for later search to find the optimal program. The full dialogue is presented in Appendix E.
@@ 5.3 Ablation studies
To justify the effectiveness of our proposed Pythonic-DSL strategy, we conduct experiments on generating either Python or DSL programs with the same 8 seeds and present the results in Table 3. The acceptance rate calculates the ratio of executable DSL programs, while the best return refers to the highest episodic return among the legal programs. Our method achieves the highest acceptance rate on 9 / 10 tasks and the highest best return on 7 / 10 tasks. The complete prompts are presented in Appendix F.
@@
Our proposed Scheduled HC is the most efficient search method. We compare various search methods in both programmatic and latent spaces in DoorKey and CleanHouse. In the learned latent space, we run CEM, CEBS, HC-$250$ (HC with a fixed population size of $250$), and our proposed Scheduled HC; in the programmatic space, we run HC-$k$$\in\{32,256,2048\}$ and our proposed Scheduled HC. The results presented in Figure 7 show that HC in the programmatic space achieves the best performance compared to latent space search. However, HC-$k$ with a fixed population size only specializes in specific tasks. To be specific, HC-$32$ for CleanHouse and HC-$2048$ for DoorKey. In contrast, our proposed Scheduled HC achieves comparatively the best performance in both tasks.
LLM provides a better initial search population. To verify the efficacy of using LLM-generated programs as search initialization for Scheduled HC, we compare it to using randomly sampled programs as initialization in DoorKey and CleanHouse. The results shown in Figure 7 suggest that initializing the search with LLM-generated programs significantly improves the sample efficiency.
@@ 5.4 Beyond initialization – a preliminary attempt on LLM revision
Given the encouraging empirical results from LLM-initialized programs, one might be curious whether LLM can keep revising and improving its programs once given feedback, e.g. reward, from the environment. For completeness, we additionally conduct studies with the following LLM-based revision:
  • Regeneration. Inspired by Chen et al. [14] and Olausson et al. [41], we instruct the LLM to re-generate non-duplicate programs with previously generated programs fed as inputs.
@@
  • Agent execution trace. Inspired by prior arts that leverage environment knowledge [55, 61] and code repair [14, 41], we feed the program with the best episodic return and its execution traces into the LLM.
  • Agent and program execution trace. Inspired by Hu et al. [26], we additionally point out the line of the program currently being executed as extra hints to the LLM.
Figure 8: LLM revision. The performance gain of LLM revision quickly saturates within the first two rounds.
We conduct the LLM revision experiments on DoorKey for five revision rounds, each round with $32$ programs with $5$ seeds. The results are presented in Figure 8. “Agent and program execution trace” performs best, indicating the importance of feeding LLM as much information as possible. Yet, the improvement quickly saturates within the first two rounds, and no significant gain is observed with more revision rounds. This indicates that only using LLM to revise the program may not be sufficient to solve the task. On the other hand, our proposed Scheduled HC can converge to an optimal program – and is “free,” i.e. no expensive API calls or “API usage has reached your notification threshold” email.
6 Conclusion
We propose an LLM-guided framework (LLM-GS), that leverages the programming skills and common sense of LLM to bootstrap the efficiency of assumption-free, random-guessing search methods. We propose a Pythonic-DSL framework to address the inability of LLMs to generate precise and grammatically correct programs. To further maximize the programs generated by LLMs, we design a search algorithm, Scheduled Hill Climbing, which can explore the programmatic search space to consistently improve the programs. The evaluation result on Karel environment shows the effectiveness of our proposed LLM-guided framework. Extensive ablation studies verify the necessities of our proposed Pythonic-DSL and Scheduled Hill Climbing.
@@ 6 Conclusion
Appendix A Schedule hill climbing detail
This Scheduler is designed for HC. $n$ is the number of programs evaluated, and $K_{start}$ and $K_{end}$ are hyper-parameters representing the initial and terminal number of neighbors to search, respectively. $r(n)$ is a sinusoidal function that smoothly increase from 00 to $1$ over the course of $N$ total programs evaluated in the environment. We provide detailed equation in Equation 2 and Equation 3
$\log_{2}{k(n)}$ $=(1-r(n))\log_{2}{K_{start}}+r(n)\log_{2}{K_{end}}\text{,}$ (2)
@@ Appendix B Karel DSL details
There are five symbols in the Karel DSL: statement, condition, action, boolean, and number. The production rules and Python converting rules are listed in Table 4. The statement counts mean total usage of converting one statement into two statements at the root of the abstract syntax tree. And the statement counts are limited to six in search of optimal programs. The program length is the total number of tokens used in string representation. The token limit is set to be forty-four. The depth in DSL is the recursive call of control flows. The depth is limited to four. These limitations are the same as the original settings in LEAPS [57]. The probabilities of the production rules are listed in Table 6. The mutation process selects one random node in the abstract syntax tree (AST). We can use the production rules to sample a sub-tree to replace the original node. The Karel syntax has many symbol-related parentheses, e.g. “i(” and “i)” which are related to the “IF” and “IFELSE”, thus LLM sometimes makes mistakes translating Python to DSL. We list common mistakes LLM made in Table 5 and how we solve these problems.
@@
Type Before After
Brackets removal DEF run m( move() m) DEF run m( move m)
Brackets separation DEF run m(move m) DEF run m( move m)
Brackets repairment DEF run m( move ) DEF run m( move m)
Redundant symbols removal DEF run m( move() m m) DEF run m( move m)

### v2

@@ 4.3 Optimizing the program policy with Scheduled Hill Climbing
Although the aforementioned techniques can already instruct the LLM to produce DSL programs that solve the target task to a certain degree, the episodic return can still be far from optimal. Since best-in-class LLMs are typically proprietary APIs, directly and iteratively fine-tuning them via gradient-based policy optimization methods is not possible. To further optimize the program generated by the LLM, we explore initializing search populations of search algorithms using LLM-generated programs. Via extensive experiments with the HC, CEM, and CEBS search algorithms on both program space and latent space whenever applicable, we discover a key improvement that is crucial when initializing search from LLM-generated programs – program-environment interaction scheduling for a more efficient allocation of the interaction budget. The intuition is that LLMs can often provide good initialization and therefore it makes sense to keep a small search budget, i.e. with a smaller population. After a while, if an optimal program is not yet found, we can gradually increase the budget and facilitate broader exploration. We design a scheduler based on this intuition:
@@
where $n$ represents the number of evaluated programs, $k(n)$ denotes a function indicating the current number of neighborhood programs based on the number of evaluated programs $n$. This function’s logarithm is a linear combination of the logarithms of two hyperparameters, $K_{start}$ and $K_{end}$, which signify the initial and terminal numbers of neighbors to search, respectively. The variable $r(n)$ governs the linear ratio, which itself is a sinusoidal function that gradually increases from 00 to $1$ throughout the evaluation of a total of $N$ programs (see Appendix E). With this scheduler, the number of neighbors $k$ is a function of the execution budget used, growing from $K_{start}$ to $K_{end}$. We apply this scheduler to Hill Climbing and call this method Scheduled Hill Climbing (Scheduled HC).
Overall, our best recipe consists of the following steps. First, we sample DSL programs for the LLM using the aforementioned techniques. Next, these programs are evaluated in the environment for the episodic return. In some easy tasks, some generated programs might have achieved perfect returns and thus are optimal programs. If no optimal program is found, we sort the programs based on the total reward decreasingly. Next, programs are selected from the sorted list, each serves as an initial program for the HC search, which is the best-performing search algorithm from our extensive experiments. HC searches $k$ neighbors at each step following the scheduler in Eq. 1. This process is repeated until either the optimal program is found or we meet the maximum program-environment interaction budget allowed ($N$).
@@
Figure 4: Efficiency in the Karel tasks. We compare our proposed LLM-guided search (LLM-GS) framework against existing methods, LEAPS, HPRL, CEBS, and HC in the Karel and Karel-Hard problem sets. The results show that our LLM-GS is significantly more efficient than these methods.
@@ 5.1 Evaluation setup
PRL tasks. We evaluate our proposed framework LLM-guided search (LLM-GS) using the Karel tasks from the two problem sets: Karel (Trivedi et al., 2021) (StairClimber, Maze, FourCorners, TopOff, Harvester, and CleanHouse) and Karel-Hard (Liu et al., 2023) (DoorKey, OneStroke, Seeder, and Snake). More task details can be found in Appendix C.
Baselines. We compare our proposed framework with the following methods: LEAPS (Trivedi et al., 2021), HPRL (Liu et al., 2023), CEBS (Carvalho et al., 2024), and Hill Climbing (Carvalho et al., 2024) (the current state-of-the-art). More details of these methods are presented in Appendix F.
Metrics. For a fair comparison of the sample efficiency, we use the number of programs evaluated to represent the program-environment interaction budget. A method is a more sample-efficient PRL algorithm if it achieves a higher average episodic return on a fixed budget. Specifically, for each task, the number of task variances is $C$. One program evaluated means a program is executed on all $C$ task variances to obtain the average return. We set the number of task variances $C=32$, the maximum number of program evaluation $N=10^{6}$, i.e., the interaction budget. Programs achieving an average return of 1.0 will be considered optimal.
Setup. We use GPT-4 (Achiam et al., 2023) (gpt-4-turbo-2024-04-09 with temperature=1.0, top_p=1.0) as our LLM module to generate the initial search population. The scheduler of our proposed Scheduled HC starts from $K_{start}=32$ to $K_{end}=2048$. We evaluate our LLM-GS and HC with $32$ random seeds and $5$ seeds for LEAPS, HPRL, and CEBS.
Data Leakage. When using modern LLMs, a key concern is whether the model has already memorized the content in the test bed, leading to biased or inflated results. We carefully rule out this possibility in Appendix G by examining the LLM release dates and the availability of optimal Karel programs, and by probing the LLM we use.
@@ 5.2 LLM-guided search significantly improves the sample efficiency
Figure 4 presents the experimental result, indicating that our proposed LLM-GS surpasses all the existing methods by a large margin on almost all the tasks. We also include a plot aggregating the performance across all ten tasks to clearly demonstrate the improvement in Appendix H (Figure 15).
On the Karel set, our framework completely solves these tasks and exhibits the best sample efficiency. Also, LLM-GS performs extraordinarily well on Seeder, OneStroke, and DoorKey, which are among the hardest of the Karel-Hard set. We highlight our DoorKey result, with an efficiency improvement from not converging at even 1M (baselines) to converging within around 50K (ours). Existing PRL algorithms struggle at DoorKey due to the fact that it is a two-stage task with a sparse reward function. Therefore, most of the baselines cannot escape the local maxima and converge around $0.5$. Curiously, our method starts at an episodic return of $\approx 0.5$ from the beginning and converges quickly without being trapped by the local maxima. To analyze, we examine one output program from LLM shown in Figure 5. The LLM clearly understands that DoorKey is a two-stage task and thus generates a corresponding two-stage structured program, which is much easier for later search to find the optimal program. The full dialogue is presented in Appendix D.
@@ 5.3 Ablation studies
To justify the effectiveness of our proposed Pythonic-DSL strategy, we conduct experiments on generating either Python or DSL programs with the same 8 seeds and present the results in Table 3. The acceptance rate calculates the ratio of executable DSL programs, while the best return refers to the highest episodic return among the legal programs. Our method achieves the highest acceptance rate on 9 / 10 tasks and the highest best return on 7 / 10 tasks. The complete prompts are presented in Appendix I.
@@
Our proposed Scheduled HC is the most efficient search method. We compare various search methods in both programmatic and latent spaces in DoorKey and CleanHouse using LLM-initialized programs. In the learned latent space, we run CEM, CEBS, HC-$250$ (HC with a fixed population size of $250$), and our proposed Scheduled HC; in the programmatic space, we run HC-$k\in\{32,256,2048\}$ and our proposed Scheduled HC. The results presented in Figure 7 show that HC in the programmatic space achieves the best performance compared to latent space search. However, HC-$k$ with a fixed population size only specializes in specific tasks. To be specific, HC-$32$ for CleanHouse and HC-$2048$ for DoorKey. In contrast, our proposed Scheduled HC achieves comparatively the best performance in both tasks.
LLM provides a better initial search population. To verify the efficacy of using LLM-generated programs as search initialization for Scheduled HC, we compare it to using randomly sampled programs as initialization in DoorKey and CleanHouse. The results shown in Figure 7 suggest that initializing the search with LLM-generated programs significantly improves the sample efficiency compared to randomly initializing population-like search methods.
5.4 Performance on novel tasks showcases the extensibility of LLM-GS
(a) PathFollow
(b) WallAvoider
Figure 8: Performance of PathFollow & WallAvoider. With only changing the task description, our proposed method surpasses the best-performing baseline HC, highlighting the extensibility of LLM-GS.
A key advantage of our proposed LLM-GS framework is its ability to adapt to new tasks using only text descriptions without requiring users to have any domain knowledge or familiarity with the DSL. To showcase this feature, we designed two novel Karel tasks, PathFollow and WallAvoider, as detailed in Section C.3. By simply replacing the task descriptions while keeping the Karel domain prompt, Python-to-DSL prompt, and hyperparameters the same, we compare our method with the best-performing baseline, HC. The results presented in Figure 8 show that by only changing the task descriptions, our LLM-GS framework still achieves improved sample efficiency and finds better policies.
@@ 5.5 Beyond initialization – a preliminary attempt on LLM revision
Given the encouraging results from LLM-initialized programs, one might be curious whether LLM can progressively revise and improve its programs once given feedback from the environment, e.g., reward. For completeness, we additionally conduct studies with the following LLM-based revision:
  • Regeneration. Inspired by Chen et al. (2024) and Olausson et al. (2024), we instruct the LLM to re-generate non-duplicate programs given its previously generated programs.
@@
  • Agent execution trace. Inspired by prior arts that leverage environment knowledge (Tang et al., 2024; Wang et al., 2024) and code repair (Chen et al., 2024; Olausson et al., 2024), we feed the program with the best episodic return and its execution traces into the LLM.
  • Agent and program execution trace. Inspired by Hu et al. (2024), we additionally point out the line of the program currently being executed as extra hints to the LLM.
Figure 9: LLM revision. The performance gain of LLM revision saturates within a few rounds.
We conduct the LLM revision experiments on DoorKey for five revision rounds, each round with $32$ programs with $5$ seeds. The results are presented in Figure 9. We can see that the improvement quickly saturates within the first two rounds, and no significant gain is observed with more revision rounds. This indicates that only using LLM to revise the program may not be sufficient to solve the task. On the other hand, our proposed Scheduled HC can converge to an optimal program – and is “free,” i.e. no expensive API calls or “API usage has reached your notification threshold” emails. The details can be found in Appendix I.
6 Discussion
We propose an LLM-guided framework (LLM-GS), that leverages the programming skills and common sense of LLM to bootstrap the efficiency of assumption-free, random-guessing search methods. We propose a Pythonic-DSL framework to address the inability of LLMs to generate precise and grammatically correct programs. To further maximize the programs generated by LLMs, we design a search algorithm, Scheduled Hill Climbing, which can explore the programmatic search space to consistently improve the programs. The experimental results on the Karel tasks show the improved sample efficiency of our proposed LLM-guided framework. Extensive ablation studies verify the effectiveness of our proposed Pythonic-DSL and Scheduled Hill Climbing. We discuss the limitations and future directions of LLM-GS in Appendix J in detail, which includes the necessity of domain experts and LLMs, etc.
@@ 6 Discussion
Appendix A Extended Related work
Sample efficiency in programmatic reinforcement learning . While deep reinforcement learning has achieved tremendous success in various domains [Silver et al., 2016, Ouyang et al., 2022, Bai et al., 2022, Wang et al., 2023b, Chen et al., 2023], it still suffers from sample inefficiency problems. Many methods make their effort to improve sample efficiency including modeling transition model [Kaiser et al., 2019], replay buffer [Mnih et al., 2015], hindsight relabeling [Andrychowicz et al., 2017], ensembling [Wiering and Van Hasselt, 2008], and importance sampling [Sutton, 2018]. Likewise, this is also the case for programmatic reinforcement learning (PRL) algorithms. The SOTA algorithm, Hill Climbing (HC), still needs to generate thousands to millions of programs to achieve optimal policies. In this work, we utilize the information on Karel’s tasks and LLM programming skills to initiate the search population of our proposed Scheduled HC to improve the sample efficiency. Incorporating other sample efficiency-improving methods into PRL algorithms is a promising research direction, but is orthogonal to our contribution.
Large language models and search-based program synthesis. Various search algorithms have been developed for program-by-example (PBE) [Gulwani, 2011, Feser et al., 2015, Polozov and Gulwani, 2015, Gulwani et al., 2017, Parisotto et al., 2017, Balog et al., 2017], whose goal is to find programs that satisfy given examples, e.g., input/output string pairs. Recent works have explored utilizing search algorithms in DSL to learn libraries [Ellis et al., 2021, Grand et al., 2024, Eberhardinger et al., 2023, Wüst et al., 2024] for PBE, imitating policies [Bastani et al., 2018, Kohler et al., 2024], and object representations [Wüst et al., 2024]. The LLM can bootstrap program synthesis in the wake stage of library learning [Grand et al., 2024, Eberhardinger et al., 2023]. It is an open question whether the LLM can bootstrap the sleep stages. We utilize the description of the Karel tasks to bootstrap our scheduled HC. Additional information may be crucial for LLM assist bootstrapping and fine-tuning. This additional information may come from the execution [Chen et al., 2024], prompt design [Wang et al., 2023a, Wei et al., 2022, Wüst et al., 2024, Grand et al., 2024], and the task description in our work.
More on search methods and structured policy representations. There are several different search methods developed for programming by example (PBE) [Gulwani, 2011, Feser et al., 2015, Polozov and Gulwani, 2015, Gulwani et al., 2017, Parisotto et al., 2017, Balog et al., 2017] or DSLs with learned libraries [Ellis et al., 2021, Grand et al., 2024, Eberhardinger et al., 2023, Wüst et al., 2024] for PBE, imitating policies [Bastani et al., 2018, Kohler et al., 2024], and object representations [Wüst et al., 2024]. Recent studies explore different representations for Reinforcement Learning policies toward interpretable and/or explainable. Together with various structured policy representations such as DSL programs [Andre and Russell, 2001, Verma et al., 2018, Silver et al., 2020], decision trees [Bastani et al., 2018, Kohler et al., 2024], state machines [Lin et al., 2024], and symbolic programs [Jiang and Luo, 2019, Qiu and Zhu, 2022, Delfosse et al., 2024]. Our method has the greatest bootstrapping ability for DSL programs but is difficult to extend to all kinds of representations. We designed our prompt by contrasting the policy representation (Karel DSL) with Python. As a result, our prompt in Section D.1 limits some of the functionalities of Python. We believe our framework is only suitable for tasks with task-solving procedures that could be described using languages. That said, it could be difficult to apply this framework to low-level control tasks, such as motor torque control [Qiu and Zhu, 2022].
@@ Appendix B Karel DSL details
There are 5 symbols in the Karel DSL: statement, condition, action, boolean, and number. The production rules and Python converting rules are listed in Table 4. The statement counts the mean total usage of converting one statement into two statements at the root of the abstract syntax tree (AST), and the statement counts are limited to 6 in search of optimal programs. The program length is the total number of tokens used in string representation and is set to 44. The depth in DSL is the recursive call of control flows, which is limited to 4. These limitations are the same as the original settings in LEAPS [Trivedi et al., 2021]. The probabilities of the production rules are listed in Table 5. The mutation process selects one random node in the AST. We use the production rules to sample a sub-tree to replace the original node. The Karel syntax has many symbol-related parentheses, e.g., “i(” and “i)” which are related to the “IF” and “IFELSE”, thus LLM sometimes makes mistakes translating Python to DSL. We list common mistakes LLM made in Table 6 and how we solve these problems.
@@
Type Before After
Brackets removal DEF run m( move() m) DEF run m( move m)
Brackets separation DEF run m(move m) DEF run m( move m)
Brackets repairment DEF run m( move ) DEF run m( move m)
Redundant symbols removal DEF run m( move() m m) DEF run m( move m)